OpenStudy (anonymous):

In a statistics class there are 18 Juniors and 10 Seniors; 6 of the seniors are female, 12 of the juniors are male. If a student is selected at random, find the probablity of selecting a) a junior or a female b) a senior or a female c)not a junior male

OpenStudy (anonymous):

a) P(junior) +P(female) b) P(senior) + P(female) c). 1 - P(male)

OpenStudy (anonymous):

So.. a) 18+6?

OpenStudy (anonymous):

Total students = senior + junior P(junior) = junior/total students

OpenStudy (anonymous):

P(A junior or a female) = (18/28) + (12/28)-(6/28) = 24/28 = 6/7 P(A senior of a female) = (10/28) + (12/28) –(6/28) = 16/28= 4/7 P(A junior or a senior) = (18/28) + (10/28) = 28/28 = 1